# A Covariant and Manifestly Projective Invariant Formulation of Thomas-Whitehead Gravity

### v1

@@ Appendix C Proofs
C.1 Proof of Eq. (3.63)
We start by calculating the trace $\mathbb{E}_{b}{}^{ab}$ via Eq. (3.50). Inserting this into Eq. (3.63) results in
$V^{a}=$ $\frac{2}{{\rm d}+1}\left\{{\bar{\nabla}}_{b}\left[g^{ab}({\mathcal{K}}+\tfrac{1}{J_{0}\kappa_{0}})-4{\mathcal{K}}^{ab}\right]-{\bar{\nabla}}_{d}(g_{bm}{\mathcal{K}}^{m(ab)d})-\lambda_{0}^{2}{\bar{{\mathcal{P}}}}_{db}K^{(ab)d}\right\}~{}~{}~{}.$ (C.1)
Using the fact that $g_{bm}{\mathcal{K}}^{mbad}$ vanishes trivially from its definition, the second term results in an additional projective Ricci term. Combining like terms and simplifying, we then have
$V^{a}=\tfrac{1}{{\rm d}+1}{\bar{\nabla}}_{b}\left[-({\rm d}-1)g^{ab}(\tfrac{1}{4J_{0}\kappa_{0}}+{\mathcal{K}})+2(2{\rm d}-3){\mathcal{K}}^{ab}\right]-\tfrac{2\lambda_{0}^{2}}{{\rm d}+1}{\bar{{\mathcal{P}}}}_{db}K^{bad}~{}~{}~{}.$ (C.2)
By inspection of Eq. (3.40), we see that the term in brackets is equivalent to $\mathbb{D}^{ab}-\lambda_{0}^{2}{\bar{\nabla}}_{c}K^{(ab)c}$. Substituting this and simplifying results in the second line of Eq. (3.63)
$V^{a}=$ $\tfrac{1}{{\rm d}+1}{\bar{\nabla}}_{b}\mathbb{D}^{ab}-\tfrac{\lambda_{0}^{2}}{{\rm d}+1}\left({\bar{\nabla}}_{b}{\bar{\nabla}}_{c}K^{(ab)c}+2{\bar{{\mathcal{P}}}}_{db}K^{bad}\right)~{}~{}~{}.$ (C.3)
C.2 Proof of Eqs. (3.65) and (3.66)
@@
$\delta\hat{\Gamma}^{c}{}_{ab}=$ $\tfrac{1}{2}\delta\left[g^{cm}(g_{m(a,b)}-g_{ab,m})\right]$ (C.4)
@@
$\delta\hat{\Gamma}^{c}{}_{ab}=$ $\tfrac{1}{2}\left[\delta g^{cm}(g_{m(a,b)}-g_{ab,m})+g^{cm}(\delta g_{m(a,b)}-\delta g_{ab,m})\right]~{}~{}~{}.$ (C.5)
@@
$\delta\hat{\Gamma}^{c}{}_{ab}=$ $\tfrac{1}{2}\big{[}\delta g^{cm}(g_{m(a,b)}-g_{ab,m})+g^{cm}(\hat{\nabla}_{(a}\delta g_{b)m}-\hat{\nabla}_{m}\delta g_{ab})$ (C.6)
$+g^{cm}\hat{\Gamma}^{p}{}_{(ab)}\delta g_{pm}+g^{cm}\hat{\Gamma}^{p}{}_{m(a}\delta g_{b)p}-g^{cm}\hat{\Gamma}^{p}{}_{m(a}\delta g_{b)p}\big{]}~{}~{}~{}.$ (C.7)
Cancelling the last two terms, expanding the Levi-Civita connection in terms of the metric in the first term on the second line, and rearranging we have
$\delta\hat{\Gamma}^{c}{}_{ab}=$ $\tfrac{1}{2}g^{cm}(\hat{\nabla}_{(a}\delta g_{b)m}-\hat{\nabla}_{m}\delta g_{ab})$ (C.8)
$+\tfrac{1}{2}\big{[}\delta g^{cm}(g_{m(a,b)}-g_{ab,m})+\delta g_{pm}g^{cm}g^{pn}(g_{n(a,b)}-g_{ab,n})\big{]}$ (C.9)
@@
$\delta\hat{\Gamma}^{c}{}_{ab}=$ $\tfrac{1}{2}g^{cm}(\hat{\nabla}_{(a}\delta g_{b)m}-\hat{\nabla}_{m}\delta g_{ab})$ (C.10)
$+\tfrac{1}{2}\big{[}\delta g^{cm}(g_{m(a,b)}-g_{ab,m})-\delta g^{cn}(g_{n(a,b)}-g_{ab,n})\big{]}~{}~{}~{}.$ (C.11)
@@
$\delta\hat{\Gamma}^{c}{}_{ab}=$ $\tfrac{1}{2}g^{cm}(\hat{\nabla}_{(a}\delta g_{b)m}-\hat{\nabla}_{m}\delta g_{ab})~{}~{}~{}.$ (C.12)
@@
$\delta\hat{\Gamma}^{m}{}_{am}=$ $-\tfrac{1}{2}g_{mn}\hat{\nabla}_{a}\delta g^{mn}~{}~{}~{}.$ (C.13)
@@
$\sqrt{g}A_{c}{}^{ab}\delta\hat{\Gamma}^{c}{}_{ab}$ (C.14)
where $A_{c}{}^{ab}=A_{c}{}^{ba}$ is an arbitrary tensor symmetric in its last two indices. Substituting Eq. (C.12) we have
$\sqrt{g}A_{c}{}^{ab}\delta\hat{\Gamma}^{c}{}_{ab}=$ $\sqrt{g}A_{c}{}^{ab}\tfrac{1}{2}g^{cm}(\hat{\nabla}_{(a}\delta g_{b)m}-\hat{\nabla}_{m}\delta g_{ab})~{}~{}~{}.$ (C.15)
@@
$\sqrt{g}A_{c}{}^{ab}\delta\hat{\Gamma}^{c}{}_{ab}=$ $\sqrt{g}\tfrac{1}{2}g^{cm}(\delta g_{ab}\hat{\nabla}_{m}A_{c}{}^{ab}-2\delta g_{bm}\hat{\nabla}_{a}A_{c}{}^{ab})~{}~{}~{}.$ (C.16)
@@
$\sqrt{g}A_{c}{}^{ab}\delta\hat{\Gamma}^{c}{}_{ab}=$ $\sqrt{g}\tfrac{1}{2}\left[\hat{\nabla}_{a}(A^{amn}-2A^{mna})\right]\delta g_{mn}~{}~{}~{}.$ (C.17)
@@
Raising indices on the variation of the metric at the cost of a minus sign and using symmetry of the metric indices to symmetrize indices on $A_{mn}{}^{a}$ yields Eq. (3.65)
$\sqrt{g}A_{c}{}^{ab}\delta\hat{\Gamma}^{c}{}_{ab}=$ $\sqrt{g}\tfrac{1}{2}\left[\hat{\nabla}_{a}(A_{(mn)}{}^{a}-A^{a}{}_{mn})\right]\delta g^{mn}~{}~{}~{}.$ (C.18)
@@
Doing the corresponding calculation for the variation of the trace contracted with an arbitrary vector $A^{a}$ quickly yields Eq. (3.66)
$\sqrt{g}A^{a}\delta\hat{\Gamma}^{m}{}_{am}=$ $\sqrt{g}A^{a}(-\tfrac{1}{2}g_{mn}\hat{\nabla}_{a}\delta g^{mn})$ (C.19)
$=$ $\sqrt{g}\tfrac{1}{2}g_{mn}(\hat{\nabla}_{a}A^{a})\delta g^{mn}$ (C.20)
$=$ $\sqrt{g}\tfrac{1}{2}g_{mn}({\bar{\nabla}}_{a}A^{a})\delta g^{mn}~{}~{}~{}.$ (C.21)
@@
The second equality holds, again, up to a total derivative. In going from the second to third lines, we have used the interchangeability of the divergences of ^$\hat{\nabla}_{a}$ and ¯${\bar{\nabla}}_{a}$, as proved in Eq. (3.22).
C.3 Proof of Eq. (3.70)
Taking the divergence of Eq. (C.3) yields
¯${\bar{\nabla}}_{a}V^{a}=\tfrac{1}{{\rm d}+1}{\bar{\nabla}}_{a}{\bar{\nabla}}_{b}\mathbb{D}^{ab}-\tfrac{\lambda_{0}^{2}}{{\rm d}+1}\left[{\bar{\nabla}}_{a}{\bar{\nabla}}_{b}{\bar{\nabla}}_{c}K^{(ab)c}+2{\bar{\nabla}}_{a}({\bar{{\mathcal{P}}}}_{db}K^{bad})\right]~{}~{}~{}.$ (C.22)
As an aside, we simplify the triple derivative as follows
¯${\bar{\nabla}}_{a}{\bar{\nabla}}_{b}{\bar{\nabla}}_{c}K^{(ab)c}=$${\bar{\nabla}}_{a}{\bar{\nabla}}_{b}{\bar{\nabla}}_{c}K^{abc}+{\bar{\nabla}}_{b}{\bar{\nabla}}_{a}{\bar{\nabla}}_{c}K^{abc}$ (C.23)
$=$ $2{\bar{\nabla}}_{a}{\bar{\nabla}}_{b}{\bar{\nabla}}_{c}K^{abc}+{\bar{R}}^{a}{}_{dba}{\bar{\nabla}}_{c}K^{dbc}+{\bar{R}}^{b}{}_{dba}{\bar{\nabla}}_{c}K^{adc}$ (C.24)
@@
where in the second line we have used the Bianchi identity for $\bar{R}^{a}{}_{bcd}$. The second and third terms cancel owing to the antisymmetry and symmetry properties ${\bar{R}}^{a}{}_{dba}=-{\bar{R}}^{a}{}_{dab}=-{\bar{R}}_{db}=-{\bar{R}}_{bd}$ and a relabeling of indices. Using the antisymmetry of $K^{abc}$ on the remaining first term, we again use the Bianchi identity to reduce this to
¯${\bar{\nabla}}_{a}{\bar{\nabla}}_{b}{\bar{\nabla}}_{c}K^{(ab)c}=$${\bar{\nabla}}_{a}({\bar{R}}^{a}{}_{dbc}K^{dbc}+{\bar{R}}^{b}{}_{dbc}K^{adc}+{\bar{R}}^{c}{}_{dbc}K^{abd})$ (C.25)
$=$ ¯${\bar{\nabla}}_{a}({\bar{R}}^{a}{}_{dbc}K^{dbc}+{\bar{R}}_{dc}K^{adc}-{\bar{R}}_{db}K^{abd})$ (C.26)
$=$ ¯${\bar{\nabla}}_{a}({\bar{R}}^{a}{}_{dbc}K^{dbc})$ (C.27)
@@
where the second and third terms have vanishing owing to the symmetric $R_{dc}$ contracted over the antisymmetric indices of $K^{adc}$. Relabeling indices and inserting into Eq. (C.22) we have
@@
¯${\bar{\nabla}}_{a}V^{a}=$ $\tfrac{1}{{\rm d}+1}{\bar{\nabla}}_{a}{\bar{\nabla}}_{b}\mathbb{D}^{ab}-\tfrac{\lambda_{0}^{2}}{{\rm d}+1}{\bar{\nabla}}_{a}\left({\bar{R}}^{a}{}_{bcd}K^{bcd}+2{\bar{{\mathcal{P}}}}_{db}K^{bad}\right)~{}~{}~{}.$ (C.28)
Using the definition of ${\mathcal{K}}^{a}{}_{bcd}$, we find the terms in the parenthesis reduce to ${\mathcal{K}}^{a}{}_{bcd}K^{bcd}$ thus we have simplified this to
¯${\bar{\nabla}}_{a}V^{a}=\tfrac{1}{{\rm d}+1}{\bar{\nabla}}_{a}{\bar{\nabla}}_{b}\mathbb{D}^{ab}-\tfrac{\lambda_{0}^{2}}{{\rm d}+1}{\bar{\nabla}}_{a}\left({\mathcal{K}}^{a}{}_{bcd}K^{bcd}\right)~{}~{}~{}.$ (3.70)

### v3

@@ Appendix C Proofs
C.1 Proof of Eq. (171)
We start by calculating the trace $\mathbb{E}_{b}{}^{ab}$ via Eq. (167). Inserting this into the definition of $V^{a}$ in Eq. (171) and simplifying results in
$V^{a}=\tfrac{-1}{{\rm d}+1}\left\{{\bar{\nabla}}_{d}\left[({\rm d}-1)\mathbb{K}^{ad}+2{\mathcal{K}}_{b}^{\ (ab)d}\right]+2\lambda_{0}^{2}{\bar{{\mathcal{P}}}}_{db}K^{(ab)d}\right\}$ (342)
with $\mathbb{K}^{ab}$ defined in Eq. (328). Using the fact that ${\mathcal{K}}_{b}^{\ bad}$ vanishes trivially from its definition, the second term results in an additional projective Ricci term. Owing to the symmetry of ${\bar{{\mathcal{P}}}}_{db}$ and antisymmetry of the last two indices of $K^{abd}$, the only non-vanishing contribution from the third term is ${\bar{{\mathcal{P}}}}_{db}K^{bad}$. Making these simplifications, relabeling indices, and combining like terms we have then
$V^{a}=$ $-\tfrac{1}{{\rm d}+1}\left\{{\bar{\nabla}}_{b}[({\rm d}-1)\mathbb{K}^{ab}+2{\mathcal{K}}^{ab}]+2\lambda_{0}^{2}{\bar{{\mathcal{P}}}}_{db}K^{bad}\right\}.$ (343)
By inspection of Eq. (158), we see that the term in brackets is equivalent to $\lambda_{0}^{2}{\bar{\nabla}}_{c}K^{(ab)c}-\mathbb{D}^{ab}$. Substituting this and simplifying results in the second line of Eq. (171)
$V^{a}=$ $\tfrac{1}{{\rm d}+1}{\bar{\nabla}}_{b}\mathbb{D}^{ab}-\tfrac{\lambda_{0}^{2}}{{\rm d}+1}\left({\bar{\nabla}}_{b}{\bar{\nabla}}_{c}K^{(ab)c}+2{\bar{{\mathcal{P}}}}_{db}K^{bad}\right).$ (344)
C.2 Proof of Eqs. (185) and (187)
@@
$\delta\hat{\Gamma}^{c}{}_{ab}=$ $\tfrac{1}{2}\delta\left[g^{cm}(g_{m(a,b)}-g_{ab,m})\right]$ (345)
@@
$\delta\hat{\Gamma}^{c}{}_{ab}=\tfrac{1}{2}\left[\delta g^{cm}(g_{m(a,b)}-g_{ab,m})+g^{cm}(\delta g_{m(a,b)}-\delta g_{ab,m})\right].$ (346)
@@
$\delta\hat{\Gamma}^{c}{}_{ab}$ $=\tfrac{1}{2}\big{[}\delta g^{cm}(g_{m(a,b)}-g_{ab,m})+g^{cm}(\hat{\nabla}_{(a}\delta g_{b)m}-\hat{\nabla}_{m}\delta g_{ab})$ (347)
$+g^{cm}\hat{\Gamma}^{p}_{\ (ab)}\delta g_{pm}+g^{cm}\hat{\Gamma}^{p}_{\ m(a}\delta g_{b)p}-g^{cm}\hat{\Gamma}^{p}_{\ m(a}\delta g_{b)p}\big{]}.$ (348)
Canceling the last two terms, expanding the Levi-Civita connection in terms of the metric in the first term on the second line, and rearranging we have
$\delta\hat{\Gamma}^{c}{}_{ab}=$ $\tfrac{1}{2}g^{cm}(\hat{\nabla}_{(a}\delta g_{b)m}-\hat{\nabla}_{m}\delta g_{ab})$ (349)
$+\tfrac{1}{2}\delta g^{cm}(g_{m(a,b)}-g_{ab,m})$ (350)
$+\tfrac{1}{2}\delta g_{pm}g^{cm}g^{pn}(g_{n(a,b)}-g_{ab,n})$ (351)
@@
$\delta\hat{\Gamma}^{c}{}_{ab}=$ $\tfrac{1}{2}g^{cm}(\hat{\nabla}_{(a}\delta g_{b)m}-\hat{\nabla}_{m}\delta g_{ab})$ (352)
$+\tfrac{1}{2}\delta g^{cm}(g_{m(a,b)}-g_{ab,m})$ (353)
$-\tfrac{1}{2}\delta g^{cn}(g_{n(a,b)}-g_{ab,n})~{}.$ (354)
@@
$\delta\hat{\Gamma}^{c}{}_{ab}=$ $\tfrac{1}{2}g^{cm}(\hat{\nabla}_{(a}\delta g_{b)m}-\hat{\nabla}_{m}\delta g_{ab})~{}.$ (355)
@@
$\delta\hat{\Gamma}^{m}{}_{am}=$ $-\tfrac{1}{2}g_{mn}\hat{\nabla}_{a}\delta g^{mn}~{}.$ (356)
@@
$\tilde{\mathbb{E}}_{c}{}^{ab}\delta\hat{\Gamma}^{c}{}_{ab}~{}.$ (357)
Substituting Eq. (355) we have
$\tilde{\mathbb{E}}_{c}{}^{ab}\delta\hat{\Gamma}^{c}{}_{ab}=$ $\tilde{\mathbb{E}}_{c}{}^{ab}\tfrac{1}{2}g^{cm}(\hat{\nabla}_{(a}\delta g_{b)m}-\hat{\nabla}_{m}\delta g_{ab}).$ (358)
@@
$\tilde{\mathbb{E}}_{c}{}^{ab}\delta\hat{\Gamma}^{c}{}_{ab}=$ $\tfrac{1}{2}g^{cm}(\delta g_{ab}\hat{\nabla}_{m}\tilde{\mathbb{E}}_{c}{}^{ab}-2\delta g_{bm}\hat{\nabla}_{a}\tilde{\mathbb{E}}_{c}{}^{ab}).$ (359)
@@
$\tilde{\mathbb{E}}_{c}{}^{ab}\delta\hat{\Gamma}^{c}{}_{ab}=$ $\tfrac{1}{2}\left[\hat{\nabla}_{a}(\tilde{\mathbb{E}}^{amn}-2\tilde{\mathbb{E}}^{mna})\right]\delta g_{mn}~{}.$ (360)
@@
Raising indices on the variation of the metric at the cost of a minus sign and using symmetry of the metric indices to symmetrize indices on $\tilde{\mathbb{E}}_{mn}{}^{a}$ yields the final result
$\tilde{\mathbb{E}}_{c}{}^{ab}\delta\hat{\Gamma}^{c}{}_{ab}=$ $\tfrac{1}{2}\delta g^{mn}\hat{\nabla}_{a}\tilde{\mathbb{E}}^{*\ \ a}_{mn}$ (185)
$\tilde{\mathbb{E}}^{*\ \ a}_{mn}\equiv$ $\tilde{\mathbb{E}}_{(mn)}^{\ \ \ \ \ a}-\tilde{\mathbb{E}}^{a}_{\ mn}$ (186)
@@
Doing the corresponding calculation for the variation of the trace $\delta\hat{\Gamma}^{m}{}_{am}$ contracted with $V^{a}$ quickly yields the final result
$V^{a}\delta\hat{\Gamma}^{m}{}_{am}=$ $V^{a}(-\tfrac{1}{2}g_{mn}\hat{\nabla}_{a}\delta g^{mn})$ (361)
$=$ $\tfrac{1}{2}\delta g^{mn}g_{mn}\hat{\nabla}_{a}V^{a}$ (362)
$=$ $\tfrac{1}{2}\delta g^{mn}g_{mn}\bar{\nabla}_{a}V^{a}~{}.$ (187)
@@
The second equality holds, again, up to a total derivative. In going from the second to third lines, we have used the interchangeability of the divergences of ^$\hat{\nabla}_{a}$ and ¯${\bar{\nabla}}_{a}$, as proved in Eq. (140).
C.3 Proof of Eq. (190)
First, we note some useful identities. Curvature as defined in Eqs. (21) and (22) can be extended to an arbitrary number of indices, and to any curvature tensor associated with its covariant derivative. Of particular use here will be the following such curvature identities
[¯$[{\bar{\nabla}}_{b},{\bar{\nabla}}_{a}]({\bar{\nabla}}_{c}K^{abc})=\bar{R}^{a}_{\ bda}{\bar{\nabla}}_{c}K^{dbc}+\bar{R}^{b}_{\ dba}{\bar{\nabla}}_{c}K^{adc}$ (363)
[¯$[{\bar{\nabla}}_{b},{\bar{\nabla}}_{c}]K^{abc}=\bar{R}^{a}_{\ dbc}K^{dbc}+\bar{R}^{b}_{\ dbc}K^{adc}+\bar{R}^{c}_{\ dbc}K^{abd}.$ (364)
We start by taking the divergence of Eq. (344), yielding
¯${\bar{\nabla}}_{a}V^{a}=\tfrac{1}{{\rm d}+1}{\bar{\nabla}}_{a}{\bar{\nabla}}_{b}\mathbb{D}^{ab}$ (365)
$\qquad-\tfrac{\lambda_{0}^{2}}{{\rm d}+1}\left[{\bar{\nabla}}_{a}{\bar{\nabla}}_{b}{\bar{\nabla}}_{c}K^{(ab)c}+2{\bar{\nabla}}_{a}({\bar{{\mathcal{P}}}}_{db}K^{bad})\right].~{}~{}~{}~{}~{}~{}$ (366)
We focus on the triple derivative, simplifying as follows
¯${\bar{\nabla}}_{a}{\bar{\nabla}}_{b}{\bar{\nabla}}_{c}K^{(ab)c}=$${\bar{\nabla}}_{a}{\bar{\nabla}}_{b}{\bar{\nabla}}_{c}K^{abc}+{\bar{\nabla}}_{b}{\bar{\nabla}}_{a}{\bar{\nabla}}_{c}K^{abc}$ (367)
$=$ $2{\bar{\nabla}}_{a}{\bar{\nabla}}_{b}{\bar{\nabla}}_{c}K^{abc}+{\bar{R}}^{a}{}_{dba}{\bar{\nabla}}_{c}K^{dbc}$ (368)
$+{\bar{R}}^{b}{}_{dba}{\bar{\nabla}}_{c}K^{adc}$ (369)
@@
where in going from the first to the second equality we have the curvature identity in Eq. (363). The second and third terms cancel owing to the antisymmetry and symmetry properties ${\bar{R}}^{a}{}_{dba}=-{\bar{R}}^{a}{}_{dab}=-{\bar{R}}_{db}=-{\bar{R}}_{bd}$ and a relabeling of indices. Using the antisymmetry of $K^{abc}$ on the remaining first term, we apply the curvature identity in Eq. (364) to reduce this to
¯${\bar{\nabla}}_{a}{\bar{\nabla}}_{b}{\bar{\nabla}}_{c}K^{(ab)c}$ =¯$={\bar{\nabla}}_{a}({\bar{R}}^{a}_{\ dbc}K^{dbc}+{\bar{R}}^{b}_{\ dbc}K^{adc}+{\bar{R}}^{c}_{\ dbc}K^{abd})$
=¯$={\bar{\nabla}}_{a}({\bar{R}}^{a}_{\ dbc}K^{dbc}+{\bar{R}}_{dc}K^{adc}-{\bar{R}}_{db}K^{abd})$
=¯$={\bar{\nabla}}_{a}({\bar{R}}^{a}_{\ dbc}K^{dbc})$ (370)
@@
where the second and third terms have vanished owing to the symmetric $\bar{R}_{dc}$ contracted over the antisymmetric indices of $K^{adc}$. Relabeling indices and inserting into Eq. (365) we have
¯${\bar{\nabla}}_{a}V^{a}=\tfrac{1}{{\rm d}+1}{\bar{\nabla}}_{a}{\bar{\nabla}}_{b}\mathbb{D}^{ab}-\tfrac{\lambda_{0}^{2}}{{\rm d}+1}{\bar{\nabla}}_{a}\left({\bar{R}}^{a}_{\ bcd}K^{bcd}+2{\bar{{\mathcal{P}}}}_{db}K^{bad}\right)$
@@
$\hskip 4.0pt=\tfrac{1}{{\rm d}+1}{\bar{\nabla}}_{a}{\bar{\nabla}}_{b}\mathbb{D}^{ab}-\tfrac{\lambda_{0}^{2}}{{\rm d}+1}{\bar{\nabla}}_{a}\left([{\bar{R}}^{a}_{\ bcd}+\delta^{a}_{\ [c}{\bar{{\mathcal{P}}}}_{d]b}]K^{bcd}\right)$ (371)
where in going from the first to last equality we have relabeled some indices, introduced a Kronecker delta symbol, and utilized the antisymmetry of the last two indices of $K^{bcd}$. Comparing with Eq. (290), we identity the terms in brackets as ${\mathcal{K}}^{a}{}_{bcd}$, thus the expression simplifies to
¯${\bar{\nabla}}_{a}V^{a}=\tfrac{1}{{\rm d}+1}{\bar{\nabla}}_{a}{\bar{\nabla}}_{b}\mathbb{D}^{ab}-\tfrac{\lambda_{0}^{2}}{{\rm d}+1}{\bar{\nabla}}_{a}\left({\mathcal{K}}^{a}{}_{bcd}K^{bcd}\right)~{}.$ (190)